# Adaptive Variation-Resilient Random Number Generator for Embedded Encryption

### v1

@@ Abstract
With a growing interest in securing user data within the internet-of-things (IoT), embedded encryption has become of paramount importance, requiring light-weight high-quality Random Number Generators (RNGs). Emerging stochastic device technologies produce random numbers from stochastic physical processes at high quality, however, their generated random number streams are adversely affected by process and supply voltage variations, which can lead to bias in the generated streams. In this work, we present an adaptive variation-resilient RNG capable of extracting unbiased encryption-grade random number streams from physically driven entropy sources, for embedded cryptography applications. As a proof of concept, we employ a stochastic magnetic tunnel junction (sMTJ) device as an entropy source. The impact of variations in the sMTJ is mitigated by employing an adaptive digitizer with an adaptive voltage reference that dynamically tracks any stochastic signal drift or deviation, leading to unbiased random bit stream generation. The generated unbiased bit streams, due to their higher entropy, then only need to undergo simplified post-processing. Statistical randomness tests based on the National Institute of Standards and Technology (NIST) test suite are conducted on bit streams obtained using simulations and FPGA entropy source emulation experiments, validating encryption-grade randomness at a significantly reduced hardware cost, and across a wide range of process-induced device variations and supply voltage fluctuations.
@@ I Introduction
Figure 1: Block Diagram of the adaptive RNG comprising (a) an entropy source (sMTJ here), (b) an adaptive digitizer with a low-pass-filter generated moving reference voltage, and (c) compact post processing.
For cryptographic applications, the encryption key is the main secret necessary for data decryption, since it is assumed that adversaries are aware of the encryption algorithm used [3, 4, 5]. Therefore, it is crucial to generate keys through random processes that make them unlikely to be guessed. Such procedures are implemented using subsystems known as Random Number Generators (RNGs), and are of two categories: True-RNGs (TRNGs) and Pseudo-RNGs (PRNGs). The latter include circuits that implement mathematical or computational algorithms that are typically used to generate random sequences; however, such sequences are deterministic and can be regenerated if the initial ‘seed’ is known, making them susceptible to prediction penetration. On the other hand, TRNGs are circuits that produce random bits based on physically random phenomena in emerging device technologies, making them more reliable for applications where security is of great concern [6, 7], such as wireless communication [8, 9] and remote sensing [10, 11, 12, 13].
Hardware TRNG implementations are based on the utilization of devices with unique properties or the exploitation of intrinsic noise present in classical electronic devices. TRNGs have become essential for cryptography-related IoT applications, as well as for probabilistic computing [14, 15, 16, 17]. The challenge in using TRNGs is that they are adversely affected by variations in the fabrication process of the device technology used [18], leading to bias in generated bit streams. Moreover, in IoT devices, which are mainly battery powered, fluctuations in the supply voltage can also adversely affect TRNGs [19, 20].
@@
Section II discusses entropy sources, Sec. III presents an adaptive TRNG with a true-random sMTJ entropy source and Sec. IV presents an adaptive RNG with a pseudo-random LFSR entropy source with experimental implementation. Finally, Sec. V concludes the paper.
@@ III.1 Adaptive Digitizer
Once a stochastic signal is generated by the entropy source, the signal needs to be converted from analog to digital, and this is the role of the digitizer. This is usually done by comparing the analog stochastic voltage signal with a reference analog voltage ($V_{REF}$) using a comparator that converts voltages above $V_{REF}$ to high output voltages (ideally $V_{DD}$: a digital 1) and voltages below $V_{REF}$ to low output voltages (ideally GND: a digital 0). Ideally, the desired output of the digitizer would be a random bit stream that is 50 % ones (1’s) and 50 % zeros (0’s) on average. Based on this, $V_{REF}$ is a critical design parameter. In our design, instead of applying a fixed $V_{REF}$ to the comparator, we design a low-pass filter (LPF) circuit that generates an adaptive moving $V_{REF}$ from the stochastic signal ($V_{stochastic}$) corresponding to its short-term average, as shown in Fig. 3 (d). This $V_{REF}$ adapts to changes in the stochastic signal, such as voltage drift or mismatch from the designed-for voltage levels due to process-induced or supply voltage variations.
The LPF is typically implemented using a resistor (R) and a capacitor (C) in a simple RC LPF configuration, as shown in Fig. 2 (d). The time constant $\tau_{LPF}$ of the RC circuit determines how quickly the reference voltage adapts to changes in $V_{stochastic}$, and is expressed as follows:
@@
The selection of $R$ and $C$ depends on how fast the system needs to adapt to fluctuations in $V_{stochastic}$. In a system employing an sMTJ entropy source, $f_{LPF}$ needs to be much lower than the average fluctuation frequency $f_{c}$ of the sMTJ, which can be characterized using the correlation time of magnetization in the sMTJ ($\tau_{c}$) [33]. This requires the following inequality to be satisfied:
@@
Careful design of the sMTJ can allow $\tau_{c}$ to be across a wide range of timescales from nano-seconds [34] to milli-seconds [31], accommodating the needs of different applications, as $\tau_{c}$ has an impact on the final throughput of the RNG.
@@ III.2 Post-Processing Circuit: The Trivium Cipher
Figure 4: Process-induced and supply-voltage variation analysis for the proposed adaptive TRNG. (a) Ranges of operation with resilience to variation in $V_{DD}$, $G_{0}$ and $TMR$. (b) Probability of 1’s and 0’s for: varying $V_{DD}$, varying $G_{0}$, and varying $TMR$, from left to right respectively. (c) Resilience range enhancement across the three parameters, $V_{DD}$, $G_{0}$ and $TMR$, normalized to the resilience range of the conventional design. (d) 2D pattern of a sample generated random bit stream using the presented adaptive TRNG.
@@ III.3 Simulation Results
To quantitatively evaluate the randomness qualities of the generated bit stream, statistical randomness tests are conducted. One commonly employed method is using the NIST test suite (NIST SP 800-22), which is a statistical test suite for random and pseudo-random number generators for cryptographic applications [36], and which will be used in the remainder of the presented study. The suite includes 16 tests, the P-value for each test is used as a measure of randomness and is assessed to determine the test’s success. A P-value of 1 indicates perfect randomness, while a P-value of 0 indicates no randomness. If the P-value exceeds 0.01, it indicates that the NIST test has been passed successfully.
@@ III.4 Adaptive TRNG Variation-Resilience Analysis
The results of this variation analysis for both the conventional and the proposed adaptive TRNGs are depicted in Fig. 4 (a) through the visually illustrated variation-resilient ranges, where all tests are passed. The detailed results are also shown in Table 1. For $V_{DD}$ variation analysis, the starting nominal value was 5 V and then the voltage was incremented in 0.5 V increments until the adaptive TRNG failed at 7.0 V, making 6.5 V as the upper limit of the $V_{DD}$ resilience range. The procedure was repeated again with 0.5 V decrements, leading to a 3.0 V lower limit of the $V_{DD}$ resilience range. The same procedure was also conducted for both $G_{0}$ and $TMR$, and the results are summarized in the visual illustration of Fig. 4 (a). Moreover, the detailed results are also summarized in Table I, with the average P-value for each test, across all scenarios where all the tests were passed, is recorded.
The results of Fig. 4 and Table 1 demonstrate the enhanced resilience of the adaptive TRNG to variations in $V_{DD}$, $G_{0}$ and $TMR$ compared to a conventional TRNG, reaching up to an enhancement in the variation-resilience range of 40 %, 75 % and 17 %, respectively, as illustrated in Fig. 4 (c). Fig. 4 (b), illustrates the percentage of 1s and 0s for six different cases of varying $V_{DD}$, $G_{0}$, and $TMR$, respectively. As is evident in all test cases, the percentage of 1s and 0s is close to an ideal 50 % each, thus confirming that the bit patterns are unbiased. Fig. 5 shows a visual illustrative scatter plot of the obtained average P-values from the NIST test suite in the presence of variations, demonstrating that they all fall well within the acceptable range, further highlighting the resilient nature of the adaptive TRNG.
@@
Variation Analysis Parameter $V_{DD}$ Variation $G_{0}$ Variation $TMR$ Variation
Frequency (Monobits) Pass 0.1721 8/8 Pass 0.6424 8/8 Pass 0.8238 8/8
Frequency within a Block Pass 0.4658 8/8 Pass 0.1161 8/8 Pass 0.2892 8/8
Runs Pass 0.4369 8/8 Pass 0.7284 8/8 Pass 0.6554 8/8
Longest Run of Ones in a Block Pass 0.5977 8/8 Pass 0.3048 8/8 Pass 0.3646 8/8
Binary Matrix Rank Pass 0.4143 8/8 Pass 0.3638 8/8 Pass 0.3338 8/8
Discrete Fourier Transform (Spectral) Pass 0.2287 8/8 Pass 0.4369 8/8 Pass 0.7255 8/8
Non-Overlapping Template Matching Pass 0.6114 8/8 Pass 0.1671 8/8 Pass 0.2771 8/8
Overlapping Template Matching Pass 0.7457 8/8 Pass 0.3901 8/8 Pass 0.3601 8/8
Maurer’s ”Universal Statistical” Pass 0.6514 8/8 Pass 0.5845 8/8 Pass 0.4523 8/8
Linear Complexity Pass 0.5402 8/8 Pass 0.7001 8/8 Pass 0.1062 8/8
Serial Pass 0.7153 8/8 Pass 0.0745 8/8 Pass 0.7624 8/8
Approximate Entropy Pass 0.2519 8/8 Pass 0.3214 8/8 Pass 0.5214 8/8
Cumulative Sums (Forward) Pass 0.3192 8/8 Pass 0.6984 8/8 Pass 0.6109 8/8
Cumulative Sums (Reverse) Pass 0.1762 8/8 Pass 0.3647 8/8 Pass 0.5274 8/8
Random Excursions Pass 0.8595 8/8 Pass 0.7491 8/8 Pass 0.5942 8/8
Random Excursions Variant Pass 0.5971 8/8 Pass 0.4941 8/8 Pass 0.6424 8/8
@@ IV Adaptive RNG with a Pseudo-Random Entropy Source
While TRNGs with physically driven naturally random entropy sources provide very high quality true-randomness desired for encryption, the advanced device technologies needed for these TRNGs are not accessible to all IoT device designers, nor are they always applicable in IoT applications that do not justify the high cost of custom designing chips with these technologies. In this section, we show how our adaptive RNG can address this issue by providing a low-cost highly accessible route to embedded encryption in IoT devices. To achieve this, we employ a low-quality pseudo-random entropy source, an LFSR, which can be viewed as an emulation of the sMTJ but with pseudo randomness. An LFSR is readily implementable on entry-level FPGA prototyping kits, and its design is a well-known text-book design. However, an LFSR on its own does not meet the randomness requirements of encryption, applications. Using our adaptive RNG system, we show how the use of the adaptive digitizer with simplified post-processing can increase the entropy of the LFSR, making it encryption-grade.
@@ IV.1 Enhancing LFSR Randomness with The Adaptive Digitizer
In order to enhance the entropy of the pseudo-random source, the digital output from the LFSR is converted to an analog signal serving as $V_{Stochastic}$, and is then fed into the adaptive digitizer that converts it back to digital. However, here the comparison is against the adaptive moving average - $V_{REF}$ of this stochastic signal, and hence the generated output bit stream is different than the original bit stream obtained from the LFSR. This output of the adaptive digitizer, will not be the same as the inputted digital bit stream from the LFSR, but will have higher entropy due to the conversion and subsequent adaptive digitization processes that infuses more noise (randomness) into the signal. This output will then undergo post-processing to produce a digital random bit stream with randomness qualities that fulfil the basic requirements of embedded cryptography applications.
Figure 5: A scatter plot as a visual illustration of the NIST Test Suite results for the variation analysis. Average P-values across the resilience range to variation in $V_{DD}$, $G_{0}$ and $TMR$ for the adaptive TRNG.

### v2

# arXiv:2507.05523v2[cs.ET] 28 Feb 2026

@@ Abstract
With a growing interest in securing user data within the internet-of-things (IoT), embedded encryption has become of paramount importance, requiring light-weight high-quality Random Number Generators (RNGs). Emerging stochastic device technologies produce random numbers from stochastic physical processes at high quality, however, their generated random number streams are adversely affected by process and supply voltage variations, which can lead to bias in the generated streams. In this work, we present an adaptive variation-resilient RNG capable of extracting unbiased encryption-grade random number streams from physically driven entropy sources, for embedded cryptography applications. The system’s key feature is its adaptive digitizer with an adaptive reference voltage. As a proof of concept, we employ a stochastic magnetic tunnel junction (sMTJ) device as an entropy source. The impact of variations in the sMTJ is mitigated by the adaptive digitizer, which generates an adaptive short-term average reference voltage that dynamically tracks any stochastic signal drift or deviation, leading to unbiased random bit stream generation. The generated bit streams, due to their higher entropy, then only need to undergo simplified post-processing. A prototype of the adaptive RNG system was experimentally implemented using discrete electronic components and an FPGA for entropy source emulation. Statistical randomness tests based on the National Institute of Standards and Technology (NIST) test suite are conducted on bit streams obtained using the simulations as well as the discrete electronic component implementation, demonstrating that the bit streams consistently pass all 16 tests of the NIST SP 800-22 test suite with a 100% pass rate. Leveraging its simplified post-processing, the adaptive RNG shows consistent operation across a wide range of throughputs from 5 to 182 Mbps. Our results validate the system’s ability to generate encryption-grade random bit streams at reduced hardware cost for IoT devices.
@@ I Introduction
Figure 1: The adaptive RNG comprising (a) an entropy source, (b) an adaptive digitizer with a low-pass-filter generated moving reference voltage, and (c) compact post processing.
For cryptographic applications, the encryption key is the main secret necessary for data decryption, since it is assumed that adversaries are aware of the encryption algorithm used [25, 45, 49]. Therefore, it is crucial to generate keys through random processes that make them unlikely to be guessed. Such procedures are implemented using subsystems known as Random Number Generators (RNGs), and are of two categories: True-RNGs (TRNGs) and Pseudo-RNGs (PRNGs). The latter include circuits that implement mathematical or computational algorithms that are typically used to generate random sequences; however, such sequences are deterministic and can be regenerated if the initial ‘seed’ is known, making them susceptible to prediction penetration. On the other hand, TRNGs are circuits that produce random bits based on physically random phenomena in emerging device technologies, making them more reliable for applications where security is of great concern [29, 51], such as wireless communication [40, 57] and remote sensing [10, 16, 20, 18, 6, 8].
Hardware TRNG implementations are based on utilization of devices with unique properties or exploitation of intrinsic noise in classical electronic devices. TRNGs have become essential for cryptography-related IoT applications [7], as well as for probabilistic computing [21, 17, 15, 5, 4, 19]. The challenge in using TRNGs is that they are adversely affected by variations in the fabrication process of the device technology used [47, 1, 3], leading to bias in generated bit streams. Moreover, in IoT devices, which are mainly battery powered, fluctuations in the supply voltage can also adversely affect TRNGs [56, 12].
@@
Section II discusses entropy sources, Sec. III presents an adaptive TRNG with a true-random sMTJ entropy source and Sec. IV presents an adaptive RNG with a pseudo-random LFSR entropy source with experimental implementation and compares the presented adaptive RNG with state-of-the-art RNG designs. Finally, Sec. V concludes the paper.
@@ III.1 Adaptive Digitizer
Once a stochastic signal is generated by the entropy source, the signal needs to be converted from analog to digital, and this is the role of the digitizer. This is usually done by comparing the analog stochastic voltage signal with a reference analog voltage ($V_{REF}$) using a comparator that converts voltages above $V_{REF}$ to high output voltages (ideally $V_{DD}$: a digital 1) and voltages below $V_{REF}$ to low output voltages (ideally GND: a digital 0). Ideally, the desired output of the digitizer would be a random bit stream that is 50 % ones (1’s) and 50 % zeros (0’s) on average. Based on this, $V_{REF}$ is a critical design parameter. In our design, instead of applying a fixed $V_{REF}$ to the comparator, we design a low-pass filter (LPF) circuit that generates an adaptive moving $V_{REF}$ from the stochastic signal ($V_{stochastic}$) corresponding to its short-term average, as shown in Fig. 3 (d).
An important novelty of our system lies in this $V_{REF}$ that adapts to changes in the stochastic signal, such as voltage drift or mismatch from the designed-for voltage levels due to process-induced or supply voltage variations, and represents a real-time short-term average of the stochastic signal itself. This is unlike other approaches in the literature that employ a tunable reference voltage that can be tuned or controlled after analysing a sample of the generated bitstreams [39]. Such approaches are usually done in the digital domain analysing a sample of the bitstream after it has been generated and then adapting the reference voltage based on the analysis for future bitstream generation. In our approach, the short-term average of the raw stochastic signal is obtained in the analog domain and in real-time, allowing real-time compensation for any stochastic signal deviation or drift, even prior to bitstream generation. The correction and adaptation of the reference voltage take place before the comparator and without needing any information about the output of the comparator, unlike the other digital approaches that are based on feedback that comes from parts of the system that are after the comparator and require information about the output of the comparator.
The LPF that generates this adaptive $V_{REF}$ can be implemented using a resistor (R) and a capacitor (C) in a simple RC LPF configuration, as shown in Fig. 2 (d). The time constant $\tau_{LPF}$ of the RC circuit determines how quickly the reference voltage adapts to changes in $V_{stochastic}$, and is expressed as follows:
@@
The selection of $R$ and $C$ depends on how fast the system needs to adapt to fluctuations in $V_{stochastic}$. In a system employing an sMTJ entropy source, $f_{LPF}$ needs to be much lower than the average fluctuation frequency $f_{c}$ of the sMTJ, which can be characterized using the autocorrelation time of magnetization in the sMTJ ($\tau_{c}$) [32]. This requires the following inequality to be satisfied:
@@
Careful design of the sMTJ can allow $\tau_{c}$ to be across a wide range of timescales from nano-seconds to milli-seconds [13, 34, 33], accommodating the needs of different applications, as $\tau_{c}$ has an impact on the final throughput of the RNG.
@@ III.2 Post-Processing Circuit: The Trivium Cipher
Figure 4: Process-induced and supply-voltage variation analysis for the proposed adaptive TRNG. (a) Ranges of operation with resilience to variation in $V_{DD}$, $G_{0}$ and TMRTMR. (b) Probability of 1’s and 0’s for: varying $V_{DD}$, varying $G_{0}$, and varying $TMR$, from left to right respectively. (c) Resilience range enhancement across the three parameters, $V_{DD}$, $G_{0}$ and $TMR$, normalized to the resilience range of the conventional design. (d) 2D pattern of a sample generated random bit stream using the presented adaptive TRNG.
@@ III.3 Simulation Results
To quantitatively evaluate the randomness qualities of the generated bit stream, statistical randomness tests are conducted. One commonly employed method is using the NIST test suite (NIST SP 800-22), which is a statistical test suite for random and pseudo-random number generators for cryptographic applications [50], and which will be used in the remainder of the presented study. The suite includes 16 tests, the P-value for each test is used as a measure of randomness and is assessed to determine the test’s success.
@@ III.4 Adaptive TRNG Variation-Resilience Analysis
The results of this variation analysis for both the conventional and the proposed adaptive TRNGs are depicted in Fig. 4 (a) through the visually illustrated variation-resilient ranges, where all tests are passed. The detailed results are also shown in Table 1. For $V_{DD}$ variation analysis, the starting nominal value was 5 V and then the voltage was incremented in 0.5 V increments until the adaptive TRNG failed at 7.0 V, making 6.5 V as the upper limit of the $V_{DD}$ resilience range. The procedure was repeated again with 0.5 V decrements, leading to a 3.0 V lower limit of the $V_{DD}$ resilience range. The same procedure was also conducted for both $G_{0}$ and $TMR$, and the results are summarized in the visual illustration of Fig. 4 (a). Moreover, the detailed results are also summarized in Table 1, with the average P-value for each test, across all scenarios where all the tests were passed, is recorded.
@@
Variation Analysis Parameter $V_{DD}$ Variation $G_{0}$ Variation TMRTMR Variation $\tau_{c}$ Variation
Test Result P-Value Pass Rate Result P-Value Pass Rate Result P-Value Pass Rate Result P-Value Pass Rate
Frequency (Monobits) Pass 0.1721 8/8 Pass 0.6424 8/8 Pass 0.8238 8/8 Pass 0.4072 2/2
Frequency within a Block Pass 0.4658 8/8 Pass 0.1161 8/8 Pass 0.2892 8/8 Pass 0.2909 2/2
Runs Pass 0.4369 8/8 Pass 0.7284 8/8 Pass 0.6554 8/8 Pass 0.5826 2/2
Longest Run of Ones in a Block Pass 0.5977 8/8 Pass 0.3048 8/8 Pass 0.3646 8/8 Pass 0.4512 2/2
Binary Matrix Rank Pass 0.4143 8/8 Pass 0.3638 8/8 Pass 0.3338 8/8 Pass 0.3891 2/2
Discrete Fourier Transform (Spectral) Pass 0.2287 8/8 Pass 0.4369 8/8 Pass 0.7255 8/8 Pass 0.3328 2/2
Non-Overlapping Template Matching Pass 0.6114 8/8 Pass 0.1671 8/8 Pass 0.2771 8/8 Pass 0.3892 2/2
Overlapping Template Matching Pass 0.7457 8/8 Pass 0.3901 8/8 Pass 0.3601 8/8 Pass 0.5679 2/2
Maurer’s ”Universal Statistical” Pass 0.6514 8/8 Pass 0.5845 8/8 Pass 0.4523 8/8 Pass 0.6179 2/2
Linear Complexity Pass 0.5402 8/8 Pass 0.7001 8/8 Pass 0.1062 8/8 Pass 0.6201 2/2
Serial Pass 0.7153 8/8 Pass 0.0745 8/8 Pass 0.7624 8/8 Pass 0.5183 2/2
Approximate Entropy Pass 0.2519 8/8 Pass 0.3214 8/8 Pass 0.5214 8/8 Pass 0.4751 2/2
Cumulative Sums (Forward) Pass 0.3192 8/8 Pass 0.6984 8/8 Pass 0.6109 8/8 Pass 0.3419 2/2
Cumulative Sums (Reverse) Pass 0.1762 8/8 Pass 0.3647 8/8 Pass 0.5274 8/8 Pass 0.4626 2/2
Random Excursions Pass 0.8595 8/8 Pass 0.7491 8/8 Pass 0.5942 8/8 Pass 0.5456 2/2
Random Excursions Variant Pass 0.5971 8/8 Pass 0.4941 8/8 Pass 0.6424 8/8 Pass 0.5947 2/2
The results of Fig. 4 and Table 1 demonstrate the enhanced resilience of the adaptive TRNG to variations in $V_{DD}$, $G_{0}$ and $TMR$ compared to a conventional TRNG, reaching up to an enhancement in the variation-resilience range of 40 %, 75 % and 17 %, respectively, as illustrated in Fig. 4 (c). Fig. 4 (b), illustrates the percentage of 1s and 0s for six different cases of varying $V_{DD}$, $G_{0}$, and $TMR$, respectively. As is evident in all test cases, the percentage of 1s and 0s is close to an ideal 50 % each, thus confirming that the bit patterns are unbiased.
[…]
Fig. 5 shows a visual illustrative scatter plot of the obtained average P-values from the NIST test suite in the presence of variations, demonstrating that they all fall well within the acceptable range, further highlighting the resilient nature of the adaptive TRNG.
@@ IV Adaptive RNG with a Pseudo-Random Entropy Source
While TRNGs with physically driven naturally random entropy sources provide very high quality true-randomness desired for encryption, the advanced device technologies needed for these TRNGs are not accessible to all IoT device designers, nor are they always applicable in IoT applications that do not justify the high cost of custom designing chips with these technologies. In this section, we show how our adaptive RNG can address this issue by providing a low-cost highly accessible route to embedded encryption in IoT devices. To achieve this, we employ a low-quality pseudo-random entropy source, an LFSR, which can be viewed as an emulation of the sMTJ but with pseudo randomness. An LFSR is readily implementable on entry-level FPGA prototyping kits, and its design is a well-known text-book design. However, an LFSR on its own does not meet the randomness requirements of encryption applications. Using our adaptive RNG system, we show how the use of the adaptive digitizer with simplified post-processing can increase the entropy of the LFSR, making it encryption-grade.
Figure 5: A scatter plot as a visual illustration of the NIST Test Suite results for the variation analysis. Average P-values across the resilience range to variation in $V_{DD}$, $G_{0}$, $TMR$ and $\tau_{c}$ for the adaptive TRNG.
@@ IV.1 Enhancing LFSR Randomness with The Adaptive Digitizer
In order to enhance the entropy of the pseudo-random source, the digital output from the LFSR is converted to an analog signal serving as $V_{Stochastic}$, and is then fed into the adaptive digitizer that converts it back to digital. However, here the comparison is against the adaptive moving average - $V_{REF}$ - of this stochastic signal, and hence the generated output bit stream is different than the original bit stream obtained from the LFSR. This output of the adaptive digitizer will not be the same as the inputted digital bit stream from the LFSR, but will have higher entropy due to the conversion and subsequent adaptive digitization processes that infuses more noise (randomness) into the signal. This output will then undergo post-processing to produce a digital random bit stream with randomness qualities that fulfil the basic requirements of embedded cryptography applications.
Figure 6: (a) Schematic diagram of system blocks and interconnections. (b) Experimental implementation of the adaptive RNG system using discrete electronic components and an FPGA board.